# Identifiability of nonlinear ODE Models with Time-Varying Parameters: the General Analytical Solution and Applications in Viral Dynamics

### v4

@@ 8.3.6 Sixth set
Chapter 9 Further simple examples
This chapter provides three simple examples. In particular, the former is a time varying system. We only study the observability of the state and the identifiability of the parameters, constant and time varying. In other words, we simply run Algorithms 1 and 4.
9.1 First Example
We consider the system characterized by the following equations:
$\left[\begin{array}[]{ll}\dot{x}_{1}&=1\\ \dot{x}_{2}&=w_{1}\\ y&=x_{1}+x_{2}t\end{array}\right.$ (9.1)
The system is directly a special case of (2.2). In particular, $m_{u}=0$, $m_{w}=1$, $p=1$, $h_{1}(x)=x_{1}+x_{2}t$,
$g^{0}=\left[\begin{array}[]{c}1\\ 0\\ \end{array}\right],~{}~{}g^{1}=\left[\begin{array}[]{c}0\\ 1\\ \end{array}\right].$ (9.2)
9.1.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 2 provides:
$\Omega=\textnormal{span}\{\nabla h_{1}\}=\textnormal{span}\{[1,t]\}.$
$\textnormal{deg}_{w}\left({\Omega}\right)=1~{}(=m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=t$ (and the unknown input reconstructability matrix from $h_{1}$ is full rank). The system is in canonical form with respect to its unknown input and the main loop is immediately interrupted. The execution of Algorithm 1 continues with Line 45. The operation at this line ($[\widetilde{h}_{1},\ldots,\widetilde{h}_{m_{w}}]=\mathcal{S}(\Sigma,~{}\Omega)$) sets
$\widetilde{h}_{1}=h_{1}=x_{1}+x_{2}t$ (9.3)
Then, Line 46 provides:
$\mu=\left[\begin{array}[]{cc}1&0\\ \accentset{\mbox{\large.}}{\mathcal{L}}_{g^{0}}\widetilde{h}_{1}&\mathcal{L}_{g^{1}}\widetilde{h}_{1}\\ \end{array}\right]=\left[\begin{array}[]{cc}1&0\\ 1+x_{2}&t\\ \end{array}\right],~{}~{}~{}~{}\nu=\left[\begin{array}[]{cc}1&0\\ -\frac{1+x_{2}}{t}&\frac{1}{t}\\ \end{array}\right]$
Finally, from $\nu$, we obtain $\widehat{g}^{0}$, and $\widehat{g}^{1}$, by using (3.4):
$\widehat{g}^{0}=g^{0}-\frac{1+x_{2}}{t}g^{1}=\left[\begin{array}[]{c}1\\ -\frac{1+x_{2}}{t}\\ \end{array}\right],~{}~{}\widehat{g}^{1}=\frac{1}{t}g^{1}=\left[\begin{array}[]{c}0\\ \frac{1}{t}\\ \end{array}\right]$
From (3.5) we obtain that, when $m_{u}=0$, the codistribution $\widetilde{\mathcal{O}}$ vanishes. The final step of Algorithm 1 is the execution of Line 47, namely:
$\mathcal{O}=\left<\left.\widehat{g}^{0},~{}\widehat{g}^{1}\right|~{}\textnormal{span}\left\{\nabla h_{1}\right\}\right>.$
Note that, .$\accentset{\mbox{\large.}}{\mathcal{L}}_{\widehat{g}^{0}}h_{1}=\mathcal{L}_{\widehat{g}^{0}}h_{1}+\frac{\partial h_{1}}{\partial t}=0$, and $\mathcal{L}_{\widehat{g}^{1}}h_{1}=1$. The algorithm in (2.8) converges at the first step ($\Omega_{1}=\Omega_{0}$), and
$\mathcal{O}=\textnormal{span}\left\{\nabla h_{1}\right\}=\textnormal{span}\{[1,t]\}.$
The rank of $\mathcal{O}$ is $1$, which is smaller than the dimension of the state. Hence, the state is not observable.
9.1.2 Identifiability analysis
Let us check the identifiability of $w_{1}$.
$\mathcal{O}^{\bot}=\textnormal{span}\left\{\left[\begin{array}[]{l}-t\\ 1\\ \end{array}\right]\right\},$ (9.4)
We have a single symmetry of the state. It is immediate to compute the symmetry of the unknown input.
In accordance with Equation (4.6), we need to compute $\xi^{0}_{1}$, $\xi^{1}_{1}$. We have:
.$\accentset{\mbox{\large.}}{\mathcal{L}}_{g^{0}}\widetilde{h}_{1}=\mathcal{L}_{g^{0}}\widetilde{h}_{1}+\frac{\partial\widetilde{h}_{1}}{\partial t}=1+x_{2},~{}~{}~{}\mathcal{L}_{g^{1}}\widetilde{h}_{1}=t.$
$\xi^{0}_{1}=1,~{}~{}\xi^{1}_{1}=0,$
and, from (4.6),
${}^{u}\chi_{1}=-\frac{1}{t}\neq 0$
Hence, $w_{1}$ is not identifiable, even locally.
9.2 Second example
We study the system characterized by the following equations:
$\left\{\begin{array}[]{ll}\dot{x}_{1}&=x_{3}\\ \dot{x}_{2}&=x_{3}w_{1}\\ \dot{x}_{3}&=w_{2}+x_{1}w_{3}\\ y&=[x_{1},~{}x_{2}].\\ \end{array}\right.$ (9.5)
It is a special case of (2.2) with: $m_{u}=0$, $m_{w}=3$, $p=2$, $h_{1}(x)=x_{1}$, $h_{2}(x)=x_{2}$,
$g^{0}=\left[\begin{array}[]{c}x_{3}\\ 0\\ 0\\ \end{array}\right],~{}g^{1}=\left[\begin{array}[]{c}0\\ x_{3}\\ 0\\ \end{array}\right],~{}g^{2}=\left[\begin{array}[]{c}0\\ 0\\ 1\\ \end{array}\right],~{}g^{3}=\left[\begin{array}[]{c}0\\ 0\\ x_{1}\\ \end{array}\right].$ (9.6)
9.2.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 2 provides:
$\Omega=\textnormal{span}\{\nabla h_{1},\nabla h_{2}\}=\textnormal{span}\{[1,0,0],~{}[0,1,0]\}.$
$\textnormal{deg}_{w}\left({\Omega}\right)=1~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{2}}h_{1}=\mathcal{L}_{g^{3}}h_{1}=\mathcal{L}_{g^{2}}h_{2}=\mathcal{L}_{g^{3}}h_{2}=0$, and $\mathcal{L}_{g^{1}}h_{2}=x_{3}$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2}$ is 1). The system is not in canonical form with respect to its unknown input. We enter the loop at lines 9-14.
Line 11 provides:
$\widetilde{h}_{1}=h_{2}=x_{2}$
Line 12 provides: ${}^{m}\mu^{0}_{0}=1$, ${}^{m}\mu^{1}_{0}=0$, ${}^{m}\mu^{0}_{1}=0$, ${}^{m}\mu^{1}_{1}=x_{3}$ and ${}^{m}\nu^{0}_{0}=1$, ${}^{m}\nu^{1}_{0}=0$, ${}^{m}\nu^{0}_{1}=0$, ${}^{m}\nu^{1}_{1}=1/x_{3}$. obtained from (3.6).
In addition, from (3.7) we obtain: ${}^{m}\widehat{g}^{0}=g^{0}$ and ${}^{m}\widehat{g}^{1}=g^{1}/x_{3}$. Line 13 computes
$\Omega=\left<\left.\left\{\begin{array}[]{ll}g^{0}&g^{1}/x_{3}\\ g^{2}&g^{3}\\ \end{array}\right\}~{}\right|\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2}\right\}\right>,$
By a direct computation, we obtain:
$\Omega=\textnormal{span}\left\{\nabla h_{1},~{}\nabla h_{2},~{}\nabla\mathcal{L}_{g^{0}}h_{1}\right\}=\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{2},~{}\nabla x_{3}\right\}.$
We have $\textnormal{deg}_{w}\left(\Omega\right)=2$. Hence, it is larger than the previous $m~{}(=1)$ but still smaller than $m_{w}~{}(=3)$. As a result, the loop at lines 9-14 is repeated again.
Line 11 provides:
$\widetilde{h}_{2}=x_{3}.$
The tensor ${}^{m}\mu$ has 9 components. In particular, the values of the new components are: ${}^{m}\mu^{0}_{2}=~{}^{m}\mu^{1}_{2}=~{}^{m}\mu^{2}_{0}=~{}^{m}\mu^{2}_{1}=0$, and ${}^{m}\mu^{2}_{2}=1$. The values of the new components of its inverse are: ${}^{m}\nu^{0}_{2}=~{}^{m}\nu^{1}_{2}=~{}^{m}\nu^{2}_{0}=~{}^{m}\nu^{2}_{1}=0$, and ${}^{m}\nu^{2}_{2}=1$.
As result, ${}^{m}\widehat{g}^{0}=~{}^{m}\nu^{0}_{0}g^{0}+~{}^{m}\nu^{0}_{1}g^{1}+~{}^{m}\nu^{0}_{2}g^{2}=g^{0}$, ${}^{m}\widehat{g}^{1}=~{}^{m}\nu^{1}_{0}g^{0}+~{}^{m}\nu^{1}_{1}g^{1}+~{}^{m}\nu^{1}_{2}g^{2}=g^{1}/x_{3}=[0,1,0]^{T}$, and ${}^{m}\widehat{g}^{2}=~{}^{m}\nu^{2}_{0}g^{0}+~{}^{m}\nu^{2}_{1}g^{1}+~{}^{m}\nu^{2}_{2}g^{2}=g^{2}$. Line 13 provides:
$\Omega=\left<\left.\left\{\begin{array}[]{lll}{}^{m}\widehat{g}^{0},&{}^{m}\widehat{g}^{1}&{}^{m}\widehat{g}^{2}\\ &g^{3}&\\ \end{array}\right\}~{}\right|\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{2},~{}\nabla x_{3}\right\}\right>=\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{2},~{}\nabla x_{3}\right\}.$
As a result, $\textnormal{deg}_{w}(\Omega)=2=m$ and the loop at Lines 9-14 is interrupted. As $m_{u}=0$, Line 21 sets $\widehat{k}^{m}=0$.
Then, the nested loop is executed. The state is augmented, by including the third UI ($w_{3}$). The vectors that characterize the dynamics of the augmented state are:
$g^{0}=\left[\begin{array}[]{c}x_{3}\\ 0\\ x_{1}w_{3}\\ 0\\ \end{array}\right],~{}g^{1}=\left[\begin{array}[]{c}0\\ x_{3}\\ 0\\ 0\\ \end{array}\right],~{}g^{2}=\left[\begin{array}[]{c}0\\ 0\\ 1\\ 0\\ \end{array}\right],~{}g^{3}=\left[\begin{array}[]{c}0\\ 0\\ 0\\ 1\\ \end{array}\right].$ (9.7)
Line 27 provides the new ${}^{m}\mu$. It only changes its $02$ component, which becomes: ${}^{m}\mu^{0}_{2}=x_{1}w_{3}$. Similarly, the new ${}^{m}\nu$ only differs for its $02$ component, which becomes: ${}^{m}\nu^{0}_{2}=-x_{1}w_{3}$. Finally, we obtain:
${}^{m}\widehat{g}^{0}=~{}^{m}\nu^{0}_{0}g^{0}+~{}^{m}\nu^{0}_{1}g^{1}+~{}^{m}\nu^{0}_{2}g^{2}=g^{0}-x_{1}w_{3}g^{2}=[x_{3},0,0,0]^{T}$, ${}^{m}\widehat{g}^{1}=~{}^{m}\nu^{1}_{0}g^{0}+~{}^{m}\nu^{1}_{1}g^{1}+~{}^{m}\nu^{1}_{2}g^{2}=g^{1}/x_{3}=[0,1,0,0]^{T}$, and ${}^{m}\widehat{g}^{2}=~{}^{m}\nu^{2}_{0}g^{0}+~{}^{m}\nu^{2}_{1}g^{1}+~{}^{m}\nu^{2}_{2}g^{2}=g^{2}=[0,0,1,0]^{T}$.
Line 32 provides $\Omega_{1}=\Omega_{0}$. As a result, the condition at Line 33 is met, the nested and the main loop are terminated.
$\mathcal{O}=\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{2},~{}\nabla x_{3}\right\}$ (9.8)
The system is not canonic.
9.2.2 Identifiability analysis
First of all, the observability codistribution of the extended state that includes $w_{3}$, is given in (9.8). Hence, the identifiability of $w_{3}$ is directly obtained by checking if $\nabla w_{3}\in\mathcal{O}$. This is not the case and, as a result, $w_{3}$ is not identifiable, even locally.
Let us check the identifiability of $w_{1}$, $w_{2}$, and $\dot{w}_{3}$. We execute Algorithm 4. First, as the system is not canonic with respect to its unknown inputs, we have symmetries of the unknown input due to the system non canonicity (${}^{c}\chi$). In particular, from (4.5). we have a single symmetry: ${}^{c}\chi=[0,0,1]^{T}$, which expresses the unidentifiability of the third UI, i.e., $\dot{w}_{3}$.
For the remaining parameters, we compute the distribution orthogonal to the observable codistribution. We have:
$\mathcal{O}^{\bot}=\textnormal{span}\left\{\left[\begin{array}[]{l}0\\ 0\\ 0\\ 1\\ \end{array}\right]\right\},$ (9.9)
We have a single symmetry of the state. It is immediate to compute the symmetry of the unknown input.
In accordance with Equation (4.6), we need to compute $\xi^{0}_{1}$, $\xi^{1}_{1}$, $\xi^{2}_{1}$ for $w_{1}$, and $\xi^{0}_{2}$, $\xi^{1}_{2}$, $\xi^{2}_{2}$ for $w_{2}$. By a direct computation, we obtain: $\xi^{0}_{1}=\xi^{1}_{1}=\xi^{2}_{1}=0$, $\xi^{0}_{2}=x_{1}$, and $\xi^{1}_{2}=\xi^{2}_{2}=0$. As a result:
${}^{u}\chi_{1}=0,~{}~{}~{}^{u}\chi_{2}=-x_{1}\neq 0.$
Therefore, only $w_{1}$ is locally identifiable.
9.3 Third example
We study the system characterized by the following equations:
$\left\{\begin{array}[]{ll}\dot{x}_{1}&=x_{3}+w_{1}+w_{2}\\ \dot{x}_{2}&=u\\ \dot{x}_{3}&=x_{1}+u\\ y&=[x_{1},~{}x_{3}].\\ \end{array}\right.$ (9.10)
It is a special case of (2.2) with: $m_{u}=1$, $m_{w}=2$, $p=2$, $h_{1}(x)=x_{1}$, $h_{2}(x)=x_{3}$,
$g^{0}=\left[\begin{array}[]{c}x_{3}\\ 0\\ x_{1}\\ \end{array}\right],~{}g^{1}=g^{2}=\left[\begin{array}[]{c}1\\ 0\\ 0\\ \end{array}\right],~{}f^{1}=\left[\begin{array}[]{c}0\\ 1\\ 1\\ \end{array}\right]$ (9.11)
9.3.1 Observability analysis
We run Algorithm 1 to obtain the observability codistribution. Line 2 provides:
$\Omega=\textnormal{span}\{\nabla h_{1},\nabla h_{2}\}=\textnormal{span}\{[1,0,0],~{}[0,0,1]\}.$
$\textnormal{deg}_{w}\left({\Omega}\right)=1~{}(<m_{w})$, as $\mathcal{L}_{g^{1}}h_{1}=\mathcal{L}_{g^{2}}h_{1}=1$, and $\mathcal{L}_{g^{1}}h_{2}=\mathcal{L}_{g^{2}}h_{2}=0$ (and the rank of the unknown input reconstructability matrix from $h_{1},~{}h_{2}$ is 1). The system is not in canonical form with respect to its unknown input. We enter the loop at lines 9-14.
Line 11 provides:
$\widetilde{h}_{1}=h_{1}=x_{1}$
Line 12 provides: ${}^{m}\mu^{0}_{0}=1$, ${}^{m}\mu^{1}_{0}=0$, ${}^{m}\mu^{0}_{1}=x_{3}$, ${}^{m}\mu^{1}_{1}=1$ and ${}^{m}\nu^{0}_{0}=1$, ${}^{m}\nu^{1}_{0}=0$, ${}^{m}\nu^{0}_{1}=-x_{3}$, ${}^{m}\nu^{1}_{1}=1$. obtained from (3.6).
In addition, from (3.7) we obtain: ${}^{m}\widehat{g}^{0}=g^{0}-x_{3}g^{1}=[0,0,x_{1}]^{T}$ and ${}^{m}\widehat{g}^{1}=g^{1}$. Line 13 computes
$\Omega=\left<\left.f^{1},~{}\left\{\begin{array}[]{ll}{}^{m}\widehat{g}^{0}&{}^{m}\widehat{g}^{1}\\ &g^{2}\\ \end{array}\right\}~{}\right|\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{3}\right\}\right>=\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{3}\right\}.$
because $\mathcal{L}_{f^{1}}x_{1}=0$, $\mathcal{L}_{f^{1}}x_{3}=1$, $\mathcal{L}_{{}^{m}\widehat{g}^{0}}x_{1}=0$, $\mathcal{L}_{{}^{m}\widehat{g}^{0}}x_{3}=x_{1}$, $\mathcal{L}_{{}^{m}\widehat{g}^{1}}x_{1}=1$, and $\mathcal{L}_{{}^{m}\widehat{g}^{1}}x_{3}=0$.
As a result, $\textnormal{deg}_{w}(\Omega)=1=m$, and the loop at Lines 9-14 is interrupted.
As $m_{u}=1>0$, Line 19 runs Algorithms 2 and 3 to determine $\widehat{k}^{m}=s^{m}_{x}+r^{m}$. In particular, by running Algorithm 2 we obtain $s^{m}_{x}=2$ and by running Algorithm 3 we obtain $r^{m}=2$. Therefore, $\widehat{k}^{m}=4$.
Then, the nested loop is executed. The state is augmented, by including the second UI ($w_{2}$). The vectors that characterize the dynamics of the augmented state are:
$g^{0}=\left[\begin{array}[]{c}x_{3}+w_{2}\\ 0\\ x_{1}\\ 0\\ \end{array}\right],~{}g^{1}=\left[\begin{array}[]{c}1\\ 0\\ 0\\ 0\\ \end{array}\right],~{}g^{2}=\left[\begin{array}[]{c}0\\ 0\\ 0\\ 1\\ \end{array}\right],~{}f=\left[\begin{array}[]{c}0\\ 1\\ 1\\ 0\\ \end{array}\right].$ (9.12)
Line 27 provides the new ${}^{m}\mu$. It only changes its $01$ component, which becomes: ${}^{m}\mu^{0}_{1}=x_{3}+w_{2}$. Similarly, the new ${}^{m}\nu$ only differs for its $01$ component, which becomes: ${}^{m}\nu^{0}_{1}=-(x_{3}+w_{2})$. Finally, we obtain:
${}^{m}\widehat{g}^{0}=~{}^{m}\nu^{0}_{0}g^{0}+~{}^{m}\nu^{0}_{1}g^{1}=g^{0}-(x_{3}+w_{2})g^{1}=[0,0,x_{1},0]^{T}$, and ${}^{m}\widehat{g}^{1}=~{}^{m}\nu^{1}_{0}g^{0}+~{}^{m}\nu^{1}_{1}g^{1}=g^{1}$.
Line 32 provides $\Omega_{1}=\Omega_{0}$.
The condition at Line 33 is not met because $k=1<\widehat{k}^{m}=4$. On the other hand, it is immediate to realize that the codistribution $\widetilde{\mathcal{O}}$ in (3.5) is empty. As a result, the convergence is achieved and the nested and the main loop are terminated.
$\mathcal{O}=\textnormal{span}\left\{\nabla x_{1},~{}\nabla x_{3}\right\}$ (9.13)
The system is not canonic.
9.3.2 Identifiability analysis
First of all, the observability codistribution of the extended state that includes $w_{2}$, is given in (9.13). Hence, the identifiability of $w_{2}$ is directly obtained by checking if $\nabla w_{2}\in\mathcal{O}$. This is not the case and, as a result, $w_{2}$ is not identifiable, even locally.
Let us check the identifiability of $w_{1}$, and $\dot{w}_{2}$. We execute Algorithm 4. First, as the system is not canonic with respect to its unknown inputs, we have symmetries of the unknown input due to the system non canonicity (${}^{c}\chi$). In particular, from (4.5). we have a single symmetry: ${}^{c}\chi=[0,1]^{T}$, which expresses the unidentifiability of the second UI, i.e., $\dot{w}_{2}$.
For the remaining parameter, we compute the distribution orthogonal to the observable codistribution. We have:
$\mathcal{O}^{\bot}=\textnormal{span}\left\{\left[\begin{array}[]{l}0\\ 1\\ 0\\ 0\\ \end{array}\right],~{}\left[\begin{array}[]{l}0\\ 0\\ 0\\ 1\\ \end{array}\right]\right\},$ (9.14)
In accordance with Equation (4.6), we need to compute $\xi^{0}_{1}$, $\xi^{1}_{1}$. By a direct computation, we obtain, for the first generator: $\xi^{0}_{1}=\xi^{1}_{1}=0$, and, consequently ${}^{u}\chi_{1}=0$. For the second generator we obtain: $\xi^{0}_{1}=1,~{}\xi^{1}_{1}=0$, and, consequently ${}^{u}\chi_{1}=-1\neq 0$. Therefore, $w_{1}$ is not locally identifiable.
Chapter 10 Conclusion
@@ 10.4 Fourth contribution
We started by simply executing Algorithm 1 and 4 to obtain the observability of the state and the identifiability of the time-varying parameters. In the first two models, the state is unobservable and the time-varying parameters unidentifiable. In the third case, all the model parameters are unidentifiable. In the case of the HIV ODE model, and in the case of the genetic toggle switch model, the results are in contrast with the current state of the art [^1] The interested reader can find a detailed explanation of the serious error made by the authors of [3] in [5]. Note that Section 6.2 of [3] contains two distinct errors. The former, which is a simple typo, was recently acknowledged by the authors (see [4]). However, in [4], the authors still concluded that the system is observable/identifiable. The very serious error committed by the authors is highlighted in Section 3.2 of [5] and lies in an incorrect use of the Inverse Function Theorem. In particular, when using this theorem, the authors did not take into account the constraints due to the dynamics of the system, as shown in Section 3.2 of [5]..
@@
We conclude by emphasizing the generality of the two algorithms (Algorithms 1 and 4) introduced in this paper. They provide, automatically, the state observability and the parameters identifiability, of any nonlinear system, independently of its complexity and type of nonlinearity and in the presence of time-varying parameters (or unknown inputs). This generality also regards the two sets of equations in (5.5) and (5.7) here introduced, which allow us to build the set of indistinguishable states and indistinguishable unknown inputs in the presence of unobservability and unidentifiability. This paper provided a first application of all these theoretical results on two viral models and several important new properties (and also an error in the state of the art) were automatically discovered. Due to the aforementioned generality, we strongly believe that our theoretical results can be very useful in many other scientific domains, basically to investigate any system that can be described by a an ODE model with time-varying parameters. This occurs in many scientific domains, ranging from the natural and applied sciences up to the social sciences. In particular, we are interested in studying the identifiability of different astrophysical models in the context of the G-dwarf problem [52]. To solve this problem, several models have been proposed since more than half century. Many of them are based on time-varying parameters. For instance, these time-varying parameters characterize the star formation rate, the presence of gas radial flows through our galaxy, and even the possibility that the Initial Mass Function (IMF) is time-varying, as we suggested long time ago, in [53]. It would be interesting to use the results introduced by this paper to investigate the identifiability of the time-varying parameters that characterize these models in order to determine which measurements are suitable to provide sufficient information that makes these parameters identifiable and, eventually, to discriminate between all the models proposed so far.
@@
We start by proving the following property:
Lemma A.1.
Let us consider the system returned by Algorithm 1 and let us suppose that $m<m_{w}$. We have:
$\mathcal{L}_{g^{m+j}}\theta=0,~{}~{}j=1,\ldots,m_{w}-m,$
for any observable function $\theta$ (i.e., for any $\theta$ such that $\nabla\theta\in\mathcal{O}$).
Proof.
We remind the reader that, when the system is not canonic, the observability codistribution is the limit codistribution of the series $\mathcal{O}^{\infty}_{k}$ defined in Section 6.2 of [2], which are computed by Algorithm 1 in this paper and denoted by $\Omega_{k}$. From Proposition 2 in [2] we know that the series of codistributions $\mathcal{O}_{k}^{\infty}$ converges and the limit codistribution is precisely $\mathcal{O}$ (see also Line 36 of Algorithm 1). Hence, the codistribution $\mathcal{O}$ is invariant under the Lie derivatives along the vector fields ${}^{m}\widehat{g}^{\alpha}_{d\infty}$, which are defined by Equation (31) in [2], for any $\alpha=0,1,\ldots,m$. As a result, for any function $\theta$, such that $\nabla\theta\in\mathcal{O}$, we obtain that $\mathcal{L}_{{}^{m}\widehat{g}^{\alpha}_{d\infty}}\theta$ is independent of $w_{m+j}$, for any $j=1,\ldots,m-m_{w}$ and for any $\alpha=0,1,\ldots,m$ (note that, for a time varying system, and for $\alpha=0$, this result actually regards .$\accentset{\mbox{\large.}}{\mathcal{L}}_{{}^{m}\widehat{g}^{0}_{d\infty}}\theta$, with .$\accentset{\mbox{\large.}}{\mathcal{L}}_{{}^{m}\widehat{g}^{0}_{d\infty}}:=\mathcal{L}_{{}^{m}\widehat{g}^{0}_{d\infty}}+\frac{\partial}{\partial t}$). Let us consider the above function when $\alpha=0$. From Equation (31) in [2], and by knowing that ${}^{m}\nu^{0}_{0}=1$, we have:
${}^{m}\widehat{g}^{0}_{d\infty}=g^{0}_{d\infty}+\sum_{j=1}^{d}~{}^{m}\nu^{0}_{j}g^{j}_{d\infty},$
with $d=m_{w}-m$. From Equations (27) and (28) in [2], we remark that, in the above expression, only $g^{0}_{d\infty}$ depends on $w_{m+j}$ ($j=1,\ldots,m_{w}-m$). On the other hand:
$\mathcal{L}_{{}^{m}\widehat{g}^{0}_{d\infty}}\theta=\mathcal{L}_{g^{0}_{d\infty}}\theta+\sum_{j=1}^{d}~{}^{m}\nu^{0}_{j}\mathcal{L}_{g^{j}_{d\infty}}\theta$
and, in this function, only the first term, $\mathcal{L}_{g^{0}_{d\infty}}\theta$, can depend on $w_{m+j}$.
From Equation (27) in [2], we have:
$\mathcal{L}_{g^{0}_{d\infty}}\theta=\mathcal{L}_{g^{0}}\theta+\sum_{j=1}^{d}(\mathcal{L}_{g^{m+j}}\theta)~{}w_{m+j}$
Since this function must be independent of $w_{m+j}$, for any $j=1,\ldots,m-m_{w}$, we obtain:
$\mathcal{L}_{g^{m+j}}\theta=0$
[…]
Now, we are ready to prove Theorem 4.2.
@@
$X:=[x^{T},~{}w_{1},\ldots,w_{m_{w}}]^{T}$ (A.1)
and we compute its observability codistribution. We denote it by $\mathcal{O}^{X}$. To obtain $\mathcal{O}^{X}$ we could run again Algorithm 1 for the extended system characterized by $X$. However, as we will see, it is unnecessary because we already have the observability codistribution $\mathcal{O}$ of the system characterized by $x$.
Starting from the expression of the vector fields that characterize the dynamics of $x$, it is immediate to obtain the expression of the vector fields that characterize the dynamics of $X$. In particular, the drift, which will be denoted by $G^{0}$, has the following expression:
$G^{0}=\left[\begin{array}[]{c}g^{0}+\sum_{j=1}^{m_{w}}g^{j}w_{j}\\ 0_{m_{w}}\\ \end{array}\right]$ (A.2)
where $0_{m_{w}}$ is the zero $m_{w}$-column vector.
We know that, for any observable function of $x$, namely a function $\theta(x)$ such that $\nabla\theta\in\mathcal{O}$, its first order Lie derivative along $G^{0}$ is observable. In other words, we have:
.$\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\theta\in\mathcal{O}^{X}$
with .$\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}:=\mathcal{L}_{G^{0}}+\frac{\partial}{\partial t}$.
@@
Algorithm 1 also provides the functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$, which are $m$ observable functions of $x$ such that the reconstructability matrix has full rank. As they are observable, $\nabla\widetilde{h}_{i}\in\mathcal{O}$, $i=1,\ldots,m$, and, consequently,
.$\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{i}\in\mathcal{O}^{X},~{}~{}~{}i=1,\ldots,m.$
In addition, as the the reconstructability matrix from them has full rank, the covectors .$\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{1},\ldots,\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{m}$ are independent and they are also independent from the generators of $\mathcal{O}$ (because the latter are independent of the unknown inputs). As a result, the $n_{o}+m$ covectors $\nabla\theta_{1},\ldots,\nabla\theta_{n_{o}},\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{1},\ldots,\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{m}$ are independent.
On the other hand, we know that the functions $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$ are observable functions, i.e., their gradient belongs to $\mathcal{O}$. From Lemma A.1, and from the structure of $G^{0}$ in (A.2), we obtain that all the functions .$\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{1},\ldots,\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{m}$ are independent of the last $m_{w}-m$ unknown inputs (i.e., the last $m_{w}-m$ entries of $X$). As a result, the dimension of $\mathcal{O}^{X}$ cannot exceed $n_{o}+m$.
Therefore:
$\mathcal{O}^{X}=\textnormal{span}\left\{\nabla\theta_{1},\ldots,\nabla\theta_{n_{o}},~{}\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{1},\ldots,\nabla\accentset{\mbox{\large.}}{\mathcal{L}}_{G^{0}}\widetilde{h}_{m}\right\}.$ (A.3)

### v5

@@ 8.3.6 Sixth set
Chapter 9 Conclusion
@@ 9.4 Fourth contribution
We started by simply executing Algorithm 1 and 5 to obtain the observability of the state and the identifiability of the time-varying parameters. In the first two models, the state is unobservable and the time-varying parameters unidentifiable. In the third case, all the model parameters are unidentifiable. In the case of the HIV ODE model, and in the case of the genetic toggle switch model, the results are in contrast with the current state of the art [^1] The interested reader can find a detailed explanation of the serious error made by the authors of [3] in [5]. Note that Section 6.2 of [3] contains two distinct errors. The former, which is a simple typo, was recently acknowledged by the authors (see [4]). However, in [4], the authors still concluded that the system is observable/identifiable. The very serious error committed by the authors is highlighted in Section 3.2 of [5] and lies in an incorrect use of the Inverse Function Theorem. In particular, when using this theorem, the authors did not take into account the constraints due to the dynamics of the system, as shown in Section 3.2 of [5]..
@@
We conclude by emphasizing the generality of the two algorithms (Algorithms 1 and 5) introduced in this paper. They provide, automatically, the state observability and the parameters identifiability, of any nonlinear system, independently of its complexity and type of nonlinearity and in the presence of time-varying parameters (or unknown inputs). This generality also regards the two sets of equations in (5.5) and (5.7) here introduced, which allow us to build the set of indistinguishable states and indistinguishable unknown inputs in the presence of unobservability and unidentifiability. This paper provided a first application of all these theoretical results on two viral models and several important new properties (and also an error in the state of the art) were automatically discovered. Due to the aforementioned generality, we strongly believe that our theoretical results can be very useful in many other scientific domains, basically to investigate any system that can be described by a an ODE model with time-varying parameters. This occurs in many scientific domains, ranging from the natural and applied sciences up to the social sciences. In particular, we are interested in studying the identifiability of different astrophysical models in the context of the G-dwarf problem [52]. To solve this problem, several models have been proposed since more than half century. Many of them are based on time-varying parameters. For instance, these time-varying parameters characterize the star formation rate, the presence of gas radial flows through our galaxy, and even the possibility that the Initial Mass Function (IMF) is time-varying, as we suggested long time ago, in [53]. It would be interesting to use the results introduced by this paper to investigate the identifiability of the time-varying parameters that characterize these models in order to determine which measurements are suitable to provide sufficient information that makes these parameters identifiable and, eventually, to discriminate between all the models proposed so far.
@@
$X:=[x^{T},~{}v_{1},\ldots,v_{m},~{}w_{m+1},\ldots,w_{m_{w}}]^{T}$ (A.1)
and we compute its observability codistribution. We denote it by $\mathcal{O}^{X}$. We remind the reader that the quantities $v_{j}$ are given in (3.9).
@@
As $m$ is the highest unknown input degree of reconstructability, we know that the dimension of $\mathcal{O}^{X}$ is $n_{o}+m$. Therefore:
$\mathcal{O}^{X}=\textnormal{span}\left\{\nabla\theta_{1},\ldots,\nabla\theta_{n_{o}},~{}\nabla v_{1},\ldots,\nabla v_{m}\right\}.$ (A.2)